OpenStudy (anonymous):

find the slope of a line perpendicular to 2y=-6x+8

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

@anonymoustwo44 I don't agree... how come ?

OpenStudy (anonymous):

slope of this line is -3 and the slope of the line perpendicular to this is the negative of the inverse of the slope of this line or in other words, m(perpendicular)=-1/m=-1/-3=1/3

OpenStudy (anonymous):

better than giving blindly answers ! (I just wanted you to explain it )

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@Kreshnik sorry :D